# Equilibrium and nonequilibrium behaviours of the Ising ferromagnetic thick cubic shell

### v1

@@ 4 Summary:
In this manuscript, we have reported the results of both equilibrium and nonequilibrium behaviour of a thick cubic Ising ferromagnetic shell. The results are obtained using Monte Carlo simulation. In the equilibrium case, the ferro-para phase transition is studied. The transition temperature (pseudo-critical temperature $T_{c}^{p}$) is obtained from the temperature which maximizes the susceptibility. The pseudo-critical temperature has been found to be dependent on the thickness ($\Delta$) of the cubic shell. Our numerical results show that the pseudo-critical temperature increases as the thickness of the cubic shell increases. For the thinnest shell (monolayer) the pseudo-critical temperature is close to the transition temperature of a two-dimensional Ising ferromagnet. On the other hand, for maximum thickness (where the thick cubic shell becomes a solid cube), the pseudo-critical temperature is approximately equal to the Monte Carlo estimate of the ferro-para transition temperature of a solid cube. In the intermediate region, the variation of the pseudo-critical temperature with the thickness of the shell is shown with a best fitted function like $T_{c}^{p}(\Delta)=a\times\tanh(b~{}\Delta)+c$, where the values of $a$, $b$ and $c$ are $a=2.98\pm 0.19$, $b=0.321\pm 0.026$ and $c=1.43\pm 0.19$ estimated for the best fit. This is an interesting result and has not been found before in the literature. We have also studied the finite size analysis. However, our results are too imprecise to identify any scaling behaviour with critical exponents.
@@
In the nonequilibrium case, we have studied the (i) relaxation behaviour and the (ii) metastable behaviour of the thick cubic shell. In the relaxation behaviour, we have allowed the system to relax from a state where all spins are up. The state of all spin-up is the equilibrium state (ground state) for $T=0$. However, an equilibrium state at any finite temperature would be achieved through the relaxation of magnetisation. The equivalent thermal condition is prepared as $T=f\times T_{c}^{P}(\Delta)$. We have used $f=1.10~{}\rm and~{}1.15$ for all values of the thickness ($\Delta$) of the shell. The exponential relaxation is observed for all values of $\Delta$. However, as the shell-thickness ($\Delta$) increases, the relaxation time $\tau_{relax}$ is found to decrease. It may be noted here that for smaller values of thickness the relaxation time decreases rapidly, in an intermediate range of the values of the thickness we have observed a plateau (where the relaxation time is almost independent of thickness) region. For higher values of thickness the relaxation time linearly decreases with the thickness. As far as we are aware, we have not encountered similar studies or results in the literature.
@@
The decay of metastable state is another example of non-equilibrium behaviour. The Ising ferromagnet shows metastability and the metastable lifetime depends on the magnitude of the applied magnetic field and it is well studied. In this paper, we have studied the dependences of metastable lifetime on the thickness of the cubic shell of Ising ferromagnet. Interestingly, the metastable lifetime also shows a non-monotonic variation with the thickness of the cubic shell. Initially, as the value of shell-thickness increases, the metastable lifetime increases and after a specific value of the shell thickness the metastable lifetime decreases. There exists a specific value of the thickness, for which the metastable lifetime becomes maximum. These results are not found in previous reports.
@@
Our study may trigger the interest of the experimentalist to study these kinds of behaviours in real ferromagnetic sample having cavity. Depending on the size of the cavity (non-magnetic), the transition temperature and the metastable lifetime can be regulated by varying the size of the cavity inside the ferromagnetic sample.

### v3

@@ 4 Summary:
In this manuscript, we have reported the results of both equilibrium and nonequilibrium behaviours of a thick cubic Ising ferromagnetic shell, investigated through Monte Carlo simulations. The results are obtained using Monte Carlo simulation. In the equilibrium case, the ferro-para phase transition is studied. The transition temperature (pseudo-critical temperature $T_{c}^{p}$) is obtained from the temperature which maximizes the susceptibility. The pseudo-critical temperature has been found to depend remarably on the thickness ($\Delta$) of the cubic shell. Our numerical results show that the pseudo-critical temperature increases as the thickness of the cubic shell increases. For the thinnest shell (monolayer) the pseudo-critical temperature is close to the transition temperature of a two-dimensional Ising ferromagnet[25]. Conversely, in the limit of maximum thickness, where the cubic shell becomes a solid cube, the pseudo-critical temperature is approximately equal to the Monte Carlo estimate of the ferro-para transition temperature of a solid cube. These results are quite obvious. In the intermediate region, the variation of the pseudo-critical temperature with the thickness of the shell is well demonstrated by the fitting function like $T_{c}^{p}(\Delta)=a\times\tanh(b~{}\Delta)+c$, where the values of $a$, $b$ and $c$ are $a=2.98\pm 0.19$, $b=0.321\pm 0.026$ and $c=1.43\pm 0.19$ estimated for the best fit. It is worth noting that our estimated critical temperatures for $\Delta=2$ and $\Delta=3$, closely match those predicted for the layered Ising ferromagnet in mean field and interfacial approximations[29]. This is an interesting result of course and has not been found before in the literature.
@@
In the nonequilibrium case, we have investigated two key aspects of the thick cubic Ising shell: (i) the relaxation behaviour and (ii) the metastable behaviour. In the relaxation behaviour, we have allowed the system to relax from a state where all spins are up. The state of all spins are up is the equilibrium state (ground state) for $T=0$. At any finite temperature, however, the system evolves toward its true equilibrium state via the relaxation of magnetisation . The equivalent thermal condition is prepared as $T=f\times T_{c}^{P}(\Delta)$. We have used $f=1.10~{}\rm and~{}1.15$ for all values of the thickness ($\Delta$) of the shell. Under these conditions, the system exhibits exponential relaxation for all values of $\Delta$. However, as the shell-thickness ($\Delta$) increases, the relaxation time $\tau_{\rm relax}$ is found to decrease. It may be noted here that for smaller values of thickness the relaxation time decreases rapidly, in an intermediate range of the shell-thickness we have observed a plateau region where the relaxation time is almost independent of shell-thickness. For larger values of shell-thickness the relaxation time linearly decreases with the thickness. As far as we are aware, we have not encountered similar studies or results in the literature.
@@
The decay of metastable state is another remarkable feature of the non-equilibrium behaviour. The Ising ferromagnet shows metastability. In this case, the well studied[21] fact is that the metastable lifetime depends on the magnitude of the applied magnetic field. In this paper, we have studied the dependences of metastable lifetime on the thickness of the thick cubic shell of Ising ferromagnet. Interestingly, the metastable lifetime also shows a non-monotonic variation with the thickness of the cubic shell. Initially, as the value of shell-thickness increases, the metastable lifetime has been found to increase. After a specific value of the shell-thickness the metastable lifetime has been found decrease. As a result, there exists a specific value of the shell-thickness, for which the metastable lifetime becomes maximum. These results are not found in any previous study.
@@
Our study may trigger the interest of the experimentalist to investigate these kinds of behaviours in real ferromagnetic sample having cavity. Depending on the size of the cavity (non-magnetic), the transition temperature and the metastable lifetime can be regulated by varying the size of the cavity inside the ferromagnetic sample.